# Geodesic completeness, cosmological bounces and inflation

### v1

@@ I Introduction
The problem of consistently embedding cosmic inflation within quantum gravity has proved a very challenging task, as illustrated for instance by the ‘eta problem’, by the difficulties to construct consistent de Sitter backgrounds in string theory and by the so-called swampland conjectures (see Cicoli et al. (2024); Palti (2019); Yamaguchi (2011); Baumann and McAllister (2015) for reviews). One might expect a bottom-up approach to the issue to be a helpful supplement in this endeavor, analogously to the study of constraints on effective field theories based on well-tested, fundamental principles Adams et al. (2006). In the context of cosmology, and assuming a sufficiently low-energy regime such that a metric description of spacetime is applicable, one principle that one might naturally wish to impose is the absence of curvature singularities.
@@
Inflation in this respect resolves the singularity problem of standard Big Bang cosmologies (provided a consistent model of reheating), by extending the spacetime at early times with a phase of quasi-exponential expansion (using cosmic time for the time coordinate) which is manifestly free of singularities. Nevertheless, even this description is not fully satisfactory, since it is known that absence of singularities does not imply geodesic completeness, i.e. the property that geodesic curves may be extended without bound into the past. Inflation precisely fails this test, as may be seen more simply in the case of exact exponential expansion: the Friedmann-Lemaître-Robertson-Walker (FLRW) metric describes only a part of the full de Sitter spacetime, and (null and time-like) geodesics cannot be extended indefinitely into the past Hawking and Ellis (2023); Vilenkin (1992). A natural question is then whether inflationary cosmologies may be extended further into the past so as to achieve geodesic completeness Aguirre and Gratton (2002); Geshnizjani et al. (2023); Yoshida and Quintin (2018); Nomura and Yoshida (2021).
@@
A famous theorem by Borde, Guth and Vilenkin (BGV) appears to forbid any such extension, at least in the context of expanding spacetimes Borde et al. (2003) (see also Borde and Vilenkin (1994); Borde (1994) for earlier work and Kinney and Stein (2022); Kinney et al. (2024); Kothawala (2021); Conroy et al. (2014); Vilenkin and Zhang (2014) for related discussions and reformulations). More in detail, the BGV theorem states that if the average expansion rate (defined below) measured by an observer is positive, then the observer’s geodesic must be past-incomplete. [^1] We say that a (null or time-like) geodesic is past- or future-complete if it may be extended indefinitely into the past or future, respectively. A complete geodesic is both past- and future-complete. We say that a spacetime is past- or future-complete if all (null or time-like) geodesics are past- or future-complete, respectively. The theorem is very general in that it does not assume the spacetime to be of the FLRW type, nor does it assume general relativity or any energy conditions. However, an interesting recent study by Easson and Lesnefsky (EL) has shown that the BGV theorem contains a subtle loophole and that, as a consequence, it may yield ‘false positives’ when diagnosing past-incompleteness Lesnefsky et al. (2023); Easson and Lesnefsky (2024a, b).
@@ II BGV theorem and its loophole
We begin by recollecting the original formulation of the BGV theorem as well as some definitions that we shall make use of in what follows. Consider a spacetime $(\mathcal{M},g)$ and let $u^{\mu}$ be a time-like vector field such that the integral curves of this field defines a future-oriented time-like geodesic congruence. Let $\mathcal{C}(\lambda)$ be a time-like or light-like geodesic curve (the ‘observer’) with tangent $v^{\mu}$ and affine parameter $\lambda$. We define the expansion rate of the congruence, or ‘generalized Hubble parameter’, with respect to the geodesic $\mathcal{C}$ as
@@
EL have proposed a simple counter-example to the BGV theorem as given in the above formulation Easson and Lesnefsky (2024a). A generalized FLRW metric with scale factor given by
@@
The trouble here is that the theorem assumes that the initial parameter $\lambda_{i}$ can be taken to be ‘strictly equal’ to $\infty$- ∞, which of course does not make sense. A complete geodesic is defined by the range $\lambda\in(-\infty,\infty)$, therefore if $H^{\mathcal{C}}_{\rm av}>0$ for all $\lambda_{i}\in(-\infty,\lambda_{f})$, then it does not follow that the geodesic is past-incomplete. Other related issues of the original BGV theorem are discussed in Ref. Lesnefsky et al. (2023).
@@
Theorem 2. Let $(\mathcal{M},g)$ be a generalized FLRW spacetime with scale factor $a(t)$. The spacetime is past null complete iff the integral $\int_{-\infty}^{t_{0}}\mathrm{d}t\,a(t)$ diverges for all $t_{0}\in\mathbb{R}$. The spacetime is past time-like complete iff the integral $\int_{-\infty}^{t_{0}}\mathrm{d}t\,\frac{a(t)}{\sqrt{a^{2}(t)+1}}$ diverges for all $t_{0}\in\mathbb{R}$.
We may verify this theorem by studying some examples. Consider first the EL model given above. It is easy to check that indeed both integrals are divergent for any $t_{0}$, confirming that this model is geodesically complete. Consider second the simplest inflationary model with $a(t)=e^{t/\alpha}$. Then $\int_{-\infty}^{t_{0}}\mathrm{d}t\,a(t)=\alpha e^{t_{0}/\alpha}$ and $\int_{-\infty}^{t_{0}}\mathrm{d}t\,\frac{a(t)}{\sqrt{a^{2}(t)+1}}=\alpha\sinh^{-1}(e^{t_{0}/\alpha})$ are both convergent and the spacetime is past-incomplete. Indeed, this spacetime is of course nothing but the expanding patch of de Sitter spacetime alluded to above.
@@ III Amendment to the BGV theorem
for all $\lambda_{0}\in(\lambda_{i},\lambda_{f})$. Then the geodesic $\mathcal{C}$ is past-incomplete. Note that $\Delta$ may in general depend on the reference parameter $\lambda_{f}$.
@@
As an alternative proof of the amended BGV theorem, but specifically in the context of generalized FLRW spacetimes, we may show that the assumptions of the theorem imply the convergence of the integrals in Theorem 2, thus implying geodesic incompleteness. Rewriting Eq. (10) we have
@@ IV Past completeness and bouncing cosmologies
Our amended BGV theorem has an interesting application in the context of generalized FLRW spacetimes. As a preliminary result we show here that past-complete cosmological spacetimes must contain at least one bounce or else be emergent (see Novello and Bergliaffa (2008); Battefeld and Peter (2015) for reviews). Here ‘bounce’ is defined in the generalized sense introduced in Ref. Easson and Lesnefsky (2024b), i.e. any time $t$ such that $\dot{a}(t)=0$, including the limiting cases where $\lim_{t\to\pm\infty}\dot{a}(t)=0$. This definition unifies several different scenarios: the usual notion of bounce, where $a(t)$ has a local minimum; a turning point corresponding to a local maximum of $a(t)$; a loitering point or phase where $\dot{a}(t)$ does not change sign but is momentarily zero; emergent cosmologies in which $a(t)$ approaches or is exactly equal to a constant in the asymptotic past. Actually, we will need here a more restricted definition of ‘emergent’, namely that $\lim_{t\to-\infty}a(t)=c$, a non-negative constant. Note that this implies that $\lim_{t\to-\infty}\dot{a}(t)=0$ only if the latter limit exists. Similarly we say that a spacetime is ‘convergent’ if $\lim_{t\to\infty}a(t)=c\geq 0$. The following lemma formalizes our result.
@@ VI Discussion
Theorem 4 is of clear importance to inflation, as it demonstrates that slow-roll models with $\epsilon\equiv-\dot{H}/H^{2}>0$ are inconsistent with geodesic completeness. An earlier phase with $\dot{H}>0$ is necessary, at least if one aims at a complete description of the pre-Big Bang epoch based on FLRW spacetime. In general relativity, $\dot{H}>0$ translates into a violation of the null energy condition (NEC). While traditionally associated with instabilities Libanov et al. (2016); Dubovsky et al. (2006); Hsu et al. (2004); Buniy and Hsu (2006); Buniy et al. (2006), there exist examples of stable models featuring violations of the NEC Creminelli et al. (2016, 2006); Nicolis et al. (2010); Creminelli et al. (2013); Nicolis and Piazza (2012); Creminelli et al. (2010); Easson et al. (2011). Intriguingly, the proposal of Creminelli et al. (2010) is precisely characterized by an emergent, past-complete FLRW metric. It would be interesting to study this question outside the context of FLRW cosmology, for instance to consider anisotropic models. Since our proposal, Theorem 3, is applicable to any spacetime, we expect it to be useful in addressing this problem. One possible such avenue would be to explore the consequences of the ‘generalized’ version of Theorem 4 for other classes of spacetimes. We leave this investigation for future work.
Acknowledgments.—This work received support from the NSFC Research Fund for International Scientists (Grant No. 12250410250).

### v2

@@ I Introduction
The problem of consistently embedding cosmic inflation within string theory has proved a very challenging task, as illustrated for instance by the ‘eta problem’, by the difficulties to construct consistent de Sitter backgrounds in supergravity and by the so-called swampland conjectures (see Cicoli et al. (2024); Palti (2019); Yamaguchi (2011); Baumann and McAllister (2015) for reviews). One might expect a bottom-up approach to the issue to be a helpful supplement in this endeavor, analogously to the study of constraints on effective field theories based on well-tested, fundamental principles Adams et al. (2006). In the context of cosmology, and assuming a sufficiently low-energy regime such that a metric description of spacetime is applicable, one principle that one might naturally wish to impose is the absence of curvature singularities.
@@
Inflation in this respect resolves the singularity problem of standard Big Bang cosmologies (provided a consistent description connecting inflation with the hot Big Bang phase), by extending the spacetime at early times with a phase of quasi-exponential expansion (using cosmic time for the time coordinate) which is manifestly free of singularities. Nevertheless, even this description is not fully satisfactory, since it is known that absence of singularities does not imply geodesic completeness, i.e. the property that geodesic curves may be extended without bound into the past. Inflation precisely fails this test, as may be seen more simply in the case of exact exponential expansion: the Friedmann-Lemaître-Robertson-Walker (FLRW) metric describes only a part of the full de Sitter spacetime, and (null and time-like) geodesics cannot be extended indefinitely into the past Hawking and Ellis (2023); Vilenkin (1992). A natural question is then whether inflationary cosmologies may be extended further into the past so as to achieve geodesic completeness Aguirre and Gratton (2002); Geshnizjani et al. (2023); Yoshida and Quintin (2018); Nomura and Yoshida (2021).
@@
A famous theorem by Borde, Guth and Vilenkin (BGV) appears to forbid any such extension, at least in the context of expanding spacetimes Borde et al. (2003) (see also Borde and Vilenkin (1994); Borde (1994) for earlier work and Kinney and Stein (2022); Kinney et al. (2024); Kothawala (2021); Conroy et al. (2014); Vilenkin and Zhang (2014) for related discussions and reformulations). More in detail, the BGV theorem states that if the average expansion rate (defined below) measured by an observer is positive, then the observer’s geodesic must be past-incomplete. [^1] We say that a (null or time-like) geodesic is past- or future-complete if it may be extended indefinitely into the past or future, respectively. A complete geodesic is both past- and future-complete. We say that a spacetime is past- or future-complete if all (null and time-like) geodesics are past- or future-complete, respectively. The theorem is very general in that it does not assume the spacetime to be of the FLRW type, nor does it assume the field equations of general relativity or any energy conditions. However, an interesting recent study by Easson and Lesnefsky (EL) has shown that the original BGV theorem contains a subtle loophole and that, as a consequence, it may yield ‘false positives’ when diagnosing past-incompleteness Lesnefsky et al. (2023); Easson and Lesnefsky (2024a, b).
@@ II BGV theorem and its loophole
We begin by recollecting the original formulation of the BGV theorem as well as some definitions that we shall make use of in what follows. Consider a spacetime $(\mathcal{M},g)$ and let $u^{\mu}$ be a time-like vector field such that the integral curves of this field define a future-oriented time-like geodesic congruence. Let $\mathcal{C}(\lambda)$ be a time-like or light-like geodesic curve (the ‘observer’) with tangent $v^{\mu}$ and affine parameter $\lambda$. We define the expansion rate of the congruence, or ‘generalized Hubble parameter’, with respect to the geodesic $\mathcal{C}$ as
@@
EL have proposed a simple example that shows the flaw in the BGV theorem as given in the above formulation Easson and Lesnefsky (2024a). A generalized FLRW metric with scale factor given by
@@
The trouble here is that the theorem assumes that the initial parameter $\lambda_{i}$ can be taken to be ‘strictly equal’ to $\infty$- ∞, which of course yields an ill-defined expression for $H^{\mathcal{C}}_{\rm av}$. A complete geodesic is defined by the range $\lambda\in(-\infty,\infty)$, therefore if $H^{\mathcal{C}}_{\rm av}>0$ for all $\lambda_{i}\in(-\infty,\lambda_{f})$, then it does not follow that the geodesic is past-incomplete. Other related issues of the original BGV theorem are discussed in Ref. Lesnefsky et al. (2023).
@@
Theorem 2. Let $(\mathcal{M},g)$ be a generalized FLRW spacetime with scale factor $a(t)$. The spacetime is past null complete iff the integral $\int_{-\infty}^{t_{0}}\mathrm{d}t\,a(t)$ diverges for all $t_{0}\in\mathbb{R}$. The spacetime is past time-like complete iff the integral $\int_{-\infty}^{t_{0}}\mathrm{d}t\,\frac{a(t)}{\sqrt{a^{2}(t)+1}}$ diverges for all $t_{0}\in\mathbb{R}$ Easson and Lesnefsky (2024b, a).
We may check this theorem by studying some examples. Consider first the EL model given above. It is easy to check that indeed both integrals are divergent for any $t_{0}$, confirming that this model is geodesically complete. Consider second the simplest inflationary model with $a(t)=e^{t/\alpha}$. Then $\int_{-\infty}^{t_{0}}\mathrm{d}t\,a(t)=\alpha e^{t_{0}/\alpha}$ and $\int_{-\infty}^{t_{0}}\mathrm{d}t\,\frac{a(t)}{\sqrt{a^{2}(t)+1}}=\alpha\sinh^{-1}(e^{t_{0}/\alpha})$ are both convergent and the spacetime is past-incomplete. Indeed, this spacetime is of course nothing but the expanding patch of de Sitter spacetime alluded to above.
@@ III Amendment to the BGV theorem
for all $\lambda_{0}\in(\lambda_{i},\lambda_{f})$. Then the geodesic $\mathcal{C}$ is past-incomplete. Note that $\Delta$ may in general depend on the reference parameter $\lambda_{f}$. [^2] After the submission of the first version of this work to the arXiv we became aware of Ref. Pavlović and Sossich (2024), where the amended version of the BGV theorem was also given. We thank Petar Pavlovic for bringing his work to our attention.
@@
As a verification of the amended BGV theorem, but specifically in the context of generalized FLRW spacetimes, we may show that the assumptions of the theorem imply the convergence of the integrals in Theorem 2, thus implying geodesic incompleteness. Rewriting Eq. (10) we have
@@ IV Past completeness and bouncing cosmologies
Our amended BGV theorem has an interesting application in the context of generalized FLRW spacetimes. As a preliminary result we show here that past-complete cosmological spacetimes must contain at least one bounce or else be emergent (see e.g. Ellis and Maartens (2004) and the reviews Novello and Bergliaffa (2008); Battefeld and Peter (2015); Brandenberger and Peter (2017)). Here ‘bounce’ is defined in the generalized sense introduced in Ref. Easson and Lesnefsky (2024b), i.e. any time $t$ such that $\dot{a}(t)=0$, including the limiting cases where $\lim_{t\to\pm\infty}\dot{a}(t)=0$. This definition unifies several different scenarios: the usual notion of bounce, where $a(t)$ has a local minimum; a turning point corresponding to a local maximum of $a(t)$; a loitering point or phase where $\dot{a}(t)$ does not change sign but is momentarily zero; emergent cosmologies in which $a(t)$ approaches or is exactly equal to a constant in the asymptotic past. Actually, we will need here a more restricted definition of ‘emergent’, namely that $\lim_{t\to-\infty}a(t)=c$, a non-negative constant. Note that this implies that $\lim_{t\to-\infty}\dot{a}(t)=0$ only if the latter limit exists. Similarly we say that a spacetime is ‘convergent’ if $\lim_{t\to\infty}a(t)=c\geq 0$. The following lemma formalizes our result.
@@ VI Discussion
Theorem 4 is of clear importance to inflation, as it demonstrates that slow-roll models with $\epsilon\equiv-\dot{H}/H^{2}>0$ are inconsistent with geodesic completeness. An earlier phase with $\dot{H}>0$ is necessary, at least if one aims at a complete description of the pre-hot Big Bang epoch based on FLRW spacetime. In general relativity, $\dot{H}>0$ typically translates into a violation of the null energy condition (NEC). Thus our theorem leads to the conclusion that any model of the early universe that respects the NEC must, within general relativity, be either singular or geodesically incomplete. This is of course in agreement with standard singularity theorems Penrose (1965); Hawking and Penrose (1970), although again our result is a priori more generally applicable.
Violations of the NEC are often associated with instabilities Libanov et al. (2016); Dubovsky et al. (2006); Hsu et al. (2004); Buniy and Hsu (2006); Buniy et al. (2006); Rubakov (2014) or, more generally, with energy densities unbounded from below in phase space Sawicki and Vikman (2013). On the other hand, at least as far as linear stability is concerned, there exist examples of stable models featuring violations of the NEC Creminelli et al. (2016, 2006); Nicolis et al. (2010); Creminelli et al. (2013); Nicolis and Piazza (2012); Creminelli et al. (2010); Easson et al. (2011). Intriguingly, the proposal of Creminelli et al. (2010) is precisely characterized by an emergent, past-complete FLRW metric where the scale factor approaches a constant in the past and evolves with increasing Hubble parameter. It would be interesting to study this question outside the context of FLRW cosmology, for instance to consider anisotropic models. Since our proposal, Theorem 3, is applicable to any spacetime, we expect it to be useful in addressing this problem. One possible such avenue would be to explore the consequences of the ‘generalized’ version of Theorem 4 for other classes of spacetimes. We leave this investigation for future work.
Acknowledgments.—We would like to thank Jun Zhang for an insightful discussion. This work received support from the NSFC Research Fund for International Scientists (Grant No. 12250410250).